OpenStudy (anonymous):

Anyone good at Algebra 1? A store allows customers to fill their own bags of candy. Troy decides he only wants chocolate-covered pretzels and gumdrops. Chocolate-covered pretzels sell for $0.89 per pound, and gumdrops sell for $0.65 per pound. Troy’s bag weighs 1.8 pounds and it cost $1.29. How many pounds of chocolate-covered pretzels and gumdrops did Troy buy? (Points : 1) 0.5 pounds of pretzels; 1.3 pounds of gumdrops 0.9 pounds of pretzels; 0.9 pounds of gumdrops 1.3 pounds of pretzels; 0.5 pounds of gumdrops 0.8 pounds of pretzels; 1 pound of gumdro

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

A=0.89*0.5+0.65*1.3=1,29 So the first answer is CORRECT :)

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Kersha has two jobs. During the day, she works as an office clerk, and in the evening she works as a cashier. Her office job pays her $12.00 per hour, and her cashier job pays her $8.25 per hour. In one week, Kersha worked 55 hours, and she earned a total of $585. If c represents the number of hours Kersha works as an office clerk, write an equation that can be used to find how many hours Kersha worked in both jobs. (Points : 1) 8.25c + 12(55 – c) = 585 12c + 8.25(585 – c) = 55 12c + 8.25(c – 5) = 585 12c + 8.25(55 – c) = 585

OpenStudy (anonymous):

I have to ask you: do you want to know the result or to be able to count it on your own?